# A Survey on Human-AI Teaming with Large Pre-Trained Models

### v1

@@ 3.1.3 UI/UX
The evolution of User Interface (UI) and User Experience (UX) design in Human-AI Joint Systems is crucial in defining the quality of collaborative experiences. Historically, UI design principles have focused on ensuring user-friendly and ethical interactions between humans and AI systems. These ethical concerns, particularly about their impact on individuals and society, have become more pressing as AI’s role in human interactions expands [45]. The complexity of such interactions demands rigorous usability testing, requiring a keen eye for detail and a profound understanding of human-computer interaction principles [44]. Traditional methodologies have highlighted the importance of clear communication and the synchronization of AI in UX evaluations, allowing AI to articulate its findings in usability testing and to determine effective modes of collaboration, whether they occur in real-time or are delayed.
@@
The recent introduction of large models has significantly transformed UI/UX design practices within Human-AI Joint Systems. Innovations such as the two-stage frameworks, with ChatIE being a prime example, have redefined zero-shot Information Extraction (IE) tasks into multi-turn question-answering sequences, marking a significant advance in the field. These models, including ChatIE, have outperformed numerous established models across various datasets [164]. Further, studies involving human subjects in varied social settings have shown that these advanced approaches not only improve the quality of interactions but also lead to more engaging and emotionally gratifying user experiences. This reflects the burgeoning capability of artificial intelligence to adopt more human-like behaviors and seamlessly integrate into social contexts [88]. This progression highlights a significant shift in the landscape of UI/UX design in Human-AI systems, where the deployment of large-scale models plays an increasingly central role in enhancing user experience and interaction fidelity.
3.2 Effective Human-AI Collaboration
@@ 3.2.1 Human Helps AI
Data annotation and preparation, as highlighted by Wei et al. [164], serve as the bedrock for training AI systems, ensuring they are equipped with accurate and contextually relevant data. This foundational work is critical for the creation of AI models that are both effective and trustworthy. The customization and refinement of AI algorithms, driven by human feedback and insights, further illustrate the central role of human expertise in AI development, enhancing the system’s accuracy and applicability [172]. Ethical considerations and contextual understanding are also paramount, with Pflanzer et al. [126] emphasizing the need for AI to operate within societal norms and comprehend complex human contexts—an aspect that purely algorithmic approaches may miss. The adaptation of AI systems to specific industry requirements showcases the importance of tailoring AI technologies to meet diverse application needs, thus enhancing their personalization and relevance [39, 93, 115].
@@
Interface design and communication strategies, as discussed by Zhang et al. [178], play a crucial role in making AI tools both accessible and user-friendly, bridging the gap between sophisticated AI outputs and practical user interaction. The continuous evolution of AI systems necessitates ongoing human intervention and guidance to stay abreast of advancements in the field [112]. In specialized domains like adaptive AI systems for driving, human behavioral data and feedback are utilized to simulate and incorporate diverse driving behaviors, demonstrating the integration of human subtleties into AI training for real-world applications [53]. Hybrid human-AI teaming scenarios, where control is dynamically allocated between humans and AI based on situational awareness, further highlight the collaboration between human judgment and AI efficiency [139].
@@
Recent advancements in large language models (LLMs) have also showcased the significant impact of human feedback on improving AI performance. Direct Preference Optimization (DPO), as introduced by Rafailov et al. [131], represents a significant leap forward in the customization of language models. DPO focuses on aligning model outputs with human preferences through a direct optimization process, bypassing the complexities and limitations associated with traditional reinforcement learning approaches. Similarly, the work of Nakano et al. [113] demonstrates how LLMs like GPT-3 can be fine-tuned using human feedback for tasks such as web-assisted question-answering. Ouyang et al. [121] further explore the benefits of fine-tuning LLMs with human feedback through the development of instructGPT models. The research of Ziegler et al. [181] complements these findings by applying reward learning, informed by human judgments, to tailor language models for specific tasks such as stylized text generation and summarization. This approach underscores the versatility of human feedback in enhancing AI models, enabling them to adapt to a wide range of stylistic preferences and content requirements. These contributions collectively underscore the critical role of humans in shaping AI development, ensuring that AI systems not only advance in capability but also align with ethical standards, societal norms, and user expectations, thereby fostering a constructive and sustainable human-AI partnership.
3.2.2 AI Helps Human
@@
Initial discussions focus on traditional AI’s role in simplifying complex tasks, like transforming zero-shot information extraction into manageable question-answering formats, aiding humans in navigating large datasets effectively [164]. Additionally, AI’s contribution to enhancing human decision-making, particularly under uncertainty, is explored through its applications in autonomous vehicles, emphasizing the need for transparency and trust in AI-human collaborations [180]. The use of ML-based systems in healthcare, such as TREWS for sepsis detection, exemplifies AI’s capacity to act as a "second pair of eyes," assisting clinicians with the voluminous clinical data for timely interventions [63]. Experimental studies on team dynamics, incorporating both humans and AI agents, reveal the efficiency gains from AI integration, where AI handles tasks suited to its capabilities, thus optimizing team performance [103]. Furthermore, innovations in human-machine collaboration, such as intelligent haptic interfaces for vehicular control, showcase improvements in safety and efficiency during the transition from automated to manual driving [97].
@@
More recently, impressive capabilities of LLMs in assisting humans, like chaining multiple LLM prompts offer a novel approach to complex task management, enhancing the human-AI collaborative experience through a transparent and controllable interface [172]. LLMs’ role in augmenting human tasks is further detailed, highlighting their utility in providing insights and suggestions for tasks that would otherwise be daunting for humans [162]. The discussion also touches on the balance required in leveraging LLMs to maintain the authenticity of human responses, underscoring the potential for high-quality but homogeneous outputs versus lower-quality but genuine human responses [157]. Interdisciplinary collaboration is advocated as a means to shape the future of LLMs and AI-human interaction, promoting a proactive and global perspective [123]. The impact of AI on human brainstorming processes, where systems like GPT-3 stimulate broader idea generation, is also examined, indicating AI’s potential to enhance cognitive processes and creativity in individual and collaborative settings [105]. The instructRL framework introduces a method where AI agents, trained through reinforcement learning, are fine-tuned with human-provided natural language instructions, demonstrating the efficacy of intuitive human-AI collaboration [65]. The application of ChatGPT in software architecture, aiding in the translation of architectural narratives into technical specifications, exemplifies AI’s role in enhancing productivity and fostering creative design processes [1]. Lastly, the examination of AI’s role in driving safety and autonomous systems, via a delegation manager that intelligently assigns control based on the situation and capabilities, illustrates the potential for AI to significantly improve operational safety and efficiency in hybrid human-AI systems [48].
3.2.3 Combining Human and AI Strengths
@@
The essence of merging the strengths of humans and AI lies in leveraging the unique capabilities of each. It involves a careful balance where each compensates for the other’s limitations, aiming for a collaboration where the overlap in errors is minimized, thus maximizing the opportunity for mutual correction and improvement. Research indicates that while explanations of AI predictions can foster blind trust, they may not always promote the appropriate reliance needed in critical domains where human oversight is crucial [7]. Optimal collaboration requires an intricate understanding of each other’s capabilities, where humans excel in cognitive skills like context recognition, anticipation of team behaviors, and corresponding action support [21].
@@
However, recognizing and integrating these diverse forms of intelligence within Human-AI Teams (HATs) poses new challenges, necessitating an in-depth understanding of AI’s capabilities and limitations by human teammates [154, 94]. Initially, mismatches in mental models of AI’s abilities can lead to errors, but over time, as humans develop a more accurate understanding, they learn to effectively complement the AI’s functions. Advanced hybrid AI approaches, including neuro-symbolic frameworks, have been introduced to model human psychological states, aiding in the prediction and adaptation to human behaviors for enhanced cooperation [154, 114, 116].
@@
Adaptation in AI behavior, aimed at enhancing collaboration, involves learning from relevant human features and modifying behavior to align with human goals, thereby fostering a more intuitive and effective partnership [107]. Incorporating computational models of emotion into mental models allows AI agents to consider affective states in decision-making, enriching the collaboration with sophisticated, goal-oriented actions [75]. Establishing a common vocabulary and shared understanding of goals, capabilities, and emotional states further solidifies the foundation of effective Human-AI collaboration [154].
@@
Recent research initiatives showcase practical applications of this synergy. AdaTest++, an augmented tool for auditing large language models (LLMs), exemplifies the collaborative effort between humans and AI in enhancing model reliability and understanding, facilitating the identification of model failures through a combined effort [132]. In the realm of mental health support, the HAILEY system demonstrates how AI can augment human empathy with analytical insights, providing feedback-driven support for more empathetic and effective peer-to-peer conversations [140]. Additionally, the innovative integration of LLMs and diffusion models for creating visual metaphors highlights the creative potential unleashed by human-AI collaboration, where human-proposed linguistic metaphors are visually realized through AI-generated imagery, bridging the gap between conceptual ideas and visual representation [20].
3.3 Compatibility of Human-AI Systems
@@ 3.3.1 Backwards Compatibility
In the realm of AI technologies, backward compatibility plays a pivotal role in ensuring the smooth integration of new or updated systems with pre-existing frameworks. This concept is crucial when embedding AI into legacy systems, as it involves careful consideration of existing tools, processes, and practices. Weisz et al. [165] delve into how generative models contribute to the modernization of applications, emphasizing the necessity for AI systems to function cohesively within older infrastructures. In the context of traditional crowd work, Chhibber et al. [25] illustrate the significance of aligning AI tools with established workflows, showcasing the potential of generative models, like Generative Adversarial Networks (GANs) and autoencoders, to augment and refine datasets in legacy systems. Additionally, Bau et al. [9] present a novel approach using GANs to adapt image priors to specific image statistics in individual photographs, thereby tackling challenges in reproducing and manipulating high-level image attributes, as evidenced in various semantic editing tasks.
@@
Discussing the broader compatibility challenges outside just the technological perspective, Wang et al. [162] talks about the transformative effects of Large Language Models (LLMs) on the job market, underscoring the necessity of integrating these advanced technologies into existing economic structures without causing major disruptions. Eloundou et al. [41] examine the impact of LLMs on the US labor market, revealing that approximately 80% of workers could experience a significant change in their tasks, with 19% potentially seeing over half of their tasks affected. This ‘exposure’ refers to a 50% increase in task efficiency without compromising quality. Furthermore, Harrer et al. [58] highlight the potential of LLMs in generative AI applications, emphasizing the need for responsible design, human oversight, and ethical considerations to prevent the spread of misinformation. They offer insights into how these models, when developed responsibly, can serve as efficient and trustworthy tools in sectors like healthcare, addressing technology, risks, and ethical issues.
3.3.2 Learning Curve and Effective Usability
@@
In Human-AI Teaming (HAT), a critical factor is the ease and efficiency of interaction between humans and AI which significantly influences the learning curve and overall usability. McNeese et al.’s [103] study talks about the team composition in HAT, revealing that AI’s ability to take a leading role in teams can enhance performance, though it often presents an initial learning curve for human participants. Lyons et al. [98] discuss the importance of intuitive interfaces in fostering a partnership, rather than a tool-based relationship, between machines and humans. A pivotal aspect in this regard is the AI’s ability to explain its decisions, especially when those decisions may seem counter-intuitive to human collaborators. Hou et al.’s [64] research highlights that trust in AI is bolstered when humans understand the rationale behind AI decisions, fostering a sense of vulnerability and reliance on AI actions. This concept is further supported by Bansal et al. [7], who find that AI explanations significantly increase human trust in the algorithm.
@@
Exploring the potential of Large Models, Wu et al. [171] also explore that breaking down the complex human tasks into multiple LLM prompts and chaining them together helps with increased human-AI interaction towards improvement. The users improvised over the new ways of interacting with the LLMs by leveraging the sub-tasks to set calibrate model expectations and making new-unit tests for each submodule of the LLM chain. Veselovsky et al.’s [157] study, however, suggests that even though prevalent usage of LLMs generating high-quality but homogeneous responses might not truly capture the real essence of human behavior, however, they can prove to be useful too. They analyze that human workers not using LLMs in their tasks produced lower quality (less essential) information than when allowed the use of LLMs. They, hence, given the rapidly evolving large model landscape, emphasize a need to constantly monitor the co-evolution of LLM and human teaming to get to some constructive contributions.
3.3.3 Evaluation the Effectiveness of Joint Human-AI Systems

### v2

@@ 3.1.3. UI/UX
The evolution of User Interface (UI) and User Experience (UX) design in Human-AI Joint Systems is crucial in defining the quality of collaborative experiences. Historically, UI design principles have focused on ensuring user-friendly and ethical interactions between humans and AI systems. These ethical concerns, particularly about their impact on individuals and society, have become more pressing as AI’s role in human interactions expands (Flathmann et al., 2021). The complexity of such interactions demands rigorous usability testing, requiring a keen eye for detail and a profound understanding of human-computer interaction principles (Fan et al., 2022). Traditional methodologies have highlighted the importance of clear communication and the synchronization of AI in UX evaluations, allowing AI to articulate its findings in usability testing and to determine effective modes of collaboration, whether they occur in real-time or are delayed.
@@
The recent introduction of large models has significantly transformed UI/UX design practices within Human-AI Joint Systems. Innovations such as the two-stage frameworks, with ChatIE being a prime example, have redefined zero-shot Information Extraction (IE) tasks into multi-turn question-answering sequences, marking a significant advance in the field. These models, including ChatIE, have outperformed numerous established models across various datasets (Wei et al., 2023). Further, studies involving human subjects in varied social settings have shown that these advanced approaches not only improve the quality of interactions but also lead to more engaging and emotionally gratifying user experiences. This reflects the burgeoning capability of artificial intelligence to adopt more human-like behaviors and seamlessly integrate into social contexts (Lemaignan et al., 2017). This progression highlights a significant shift in the landscape of UI/UX design in Human-AI systems, where the deployment of large-scale models plays an increasingly central role in enhancing user experience and interaction fidelity.
3.2. Effective Human-AI Collaboration
@@ 3.2.1. Human Helps AI
Data annotation and preparation, as highlighted by Wei et al. (Wei et al., 2023), serve as the bedrock for training AI systems, ensuring they are equipped with accurate and contextually relevant data. The customization and refinement of AI algorithms, driven by human feedback and insights, further illustrate the central role of human expertise in AI development, enhancing the system’s accuracy and applicability (Wu et al., 2022). Ethical considerations and contextual understanding are also paramount, with Pflanzer et al. (Pflanzer et al., 2022) emphasizing the need for AI to operate within societal norms and comprehend complex human contexts—an aspect that purely algorithmic approaches may miss. The adaptation of AI systems to specific industry requirements showcases the importance of tailoring AI technologies to meet diverse application needs, thus enhancing their personalization and relevance (Dubey et al., 2020; Liu et al., 2023a; Ngo et al., 2023).
@@
Interface design and communication strategies, as discussed by Zhang et al. (Zhang et al., 2023), play a crucial role in making AI tools both accessible and user-friendly, bridging the gap between sophisticated AI outputs and practical user interaction. The continuous evolution of AI systems necessitates ongoing human intervention and guidance to stay abreast of advancements in the field (Munyaka et al., 2023). In specialized domains like adaptive AI systems for driving, human behavioral data and feedback are utilized to simulate and incorporate diverse driving behaviors, demonstrating the integration of human subtleties into AI training for real-world applications (Gopinath et al., 2022). Hybrid human-AI teaming scenarios, where control is dynamically allocated between humans and AI based on situational awareness, further highlight the collaboration between human judgment and AI efficiency (Sharifi Noorian et al., 2022).
@@
Recent advancements in large language models (LLMs) have also showcased the significant impact of human feedback on improving AI performance. Direct Preference Optimization (DPO), as introduced by Rafailov et al. (Rafailov et al., 2023b), represents a significant leap forward in the customization of language models. DPO focuses on aligning model outputs with human preferences through a direct optimization process, bypassing the complexities and limitations associated with traditional reinforcement learning approaches. Similarly, the work of Nakano et al. (Nakano et al., 2021) demonstrates how LLMs like GPT-3 can be fine-tuned using human feedback for tasks such as web-assisted question-answering. Furthermore, the instructRL framework introduces a method where AI agents, trained through reinforcement learning, are fine-tuned with human-provided natural language instructions, demonstrating the efficacy of intuitive human-AI collaboration (Hu and Sadigh, 2023). On the similar lines, Ouyang et al. (Ouyang et al., 2022) explore the benefits of fine-tuning LLMs with human feedback through the development of instructGPT models. The research of Ziegler et al. (Ziegler et al., 2019) complements these findings by applying reward learning, informed by human judgments, to tailor language models for specific tasks such as stylized text generation and summarization. This approach highlights the versatility of human feedback in enhancing AI models, enabling them to adapt to a wide range of stylistic preferences and content requirements.
3.2.2. AI Helps Human
@@
Initial discussions focus on traditional AI’s role in simplifying complex tasks, like transforming zero-shot information extraction into manageable question-answering formats, aiding humans in navigating large datasets effectively (Wei et al., 2023). Additionally, AI’s contribution to enhancing human decision-making, particularly under uncertainty, is explored through its applications in autonomous vehicles, emphasizing the need for transparency and trust in AI-human collaborations (Zhou and Chen, 2019). The use of ML-based systems in healthcare, such as TREWS for sepsis detection, exemplifies AI’s capacity to act as a "second pair of eyes," assisting clinicians with the voluminous clinical data for timely interventions (Henry et al., 2022). Experimental studies on team dynamics, incorporating both humans and AI agents, reveal the efficiency gains from AI integration, where AI handles tasks suited to its capabilities, thus optimizing team performance (McNeese et al., 2021). Furthermore, innovations in human-machine collaboration, such as intelligent haptic interfaces for vehicular control, showcase improvements in safety and efficiency during the transition from automated to manual driving (Lv et al., 2021).
@@
More recently, impressive capabilities of LLMs in assisting humans, like chaining multiple LLM prompts offer a novel approach to complex task management, enhancing the human-AI collaborative experience through a transparent and controllable interface (Wu et al., 2022). LLMs’ role in augmenting, rather than replacing, human tasks is further detailed in Wang’s work (Wang, 2023), highlighting their utility in providing insights and suggestions for tasks that would otherwise be daunting for humans. Further, Veselovsky et al. (Veselovsky et al., 2023) touch on the balance required in leveraging LLMs to maintain the authenticity of human responses, underscoring the potential for high-quality but homogeneous outputs versus lower-quality but genuine human responses. Interdisciplinary collaboration is advocated as a means to shape the future of LLMs and AI-human interaction, promoting a proactive and global perspective (Pal, 2023). The impact of AI on human brainstorming processes, where systems like GPT-3 stimulate broader idea generation, is also examined, indicating AI’s potential to enhance cognitive processes and creativity in individual and collaborative settings (Memmert and Tavanapour, 2023). The application of ChatGPT in software architecture, aiding in the translation of architectural narratives into technical specifications, exemplifies AI’s role in enhancing productivity and fostering creative design processes (Ahmad et al., 2023). Furthermore, the examination of AI’s role in driving safety and autonomous systems, via a delegation manager that intelligently assigns control based on the situation and capabilities, illustrates the potential for AI to significantly improve operational safety and efficiency in hybrid human-AI systems (Fuchs et al., 2023).
3.2.3. Combining Human and AI Strengths
@@
The essence of merging the strengths of humans and AI lies in leveraging the unique capabilities of each. It involves a careful balance where each compensates for the other’s limitations, aiming for a collaboration where the overlap in errors is minimized, thus maximizing the opportunity for mutual correction and improvement. Research indicates that while explanations of AI predictions can foster blind trust, they may not always promote the appropriate reliance needed in critical domains where human oversight is crucial (Bansal et al., 2021). Optimal collaboration requires an intricate understanding of each other’s capabilities, where humans excel in cognitive skills like context recognition, anticipation of team behaviors, and corresponding action support (Chakraborti et al., 2017).
@@
However, recognizing and integrating these diverse forms of intelligence within Human-AI Teams (HATs) poses new challenges, necessitating an in-depth understanding of AI’s capabilities and limitations by human teammates (van den Bosch et al., 2019; Liu et al., 2023b). Initially, mismatches in mental models of AI’s abilities can lead to errors, but over time, as humans develop a more accurate understanding, they learn to effectively complement the AI’s functions. Advanced hybrid AI approaches, including neuro-symbolic frameworks, have been introduced to model human psychological states, aiding in the prediction and adaptation to human behaviors for enhanced cooperation (van den Bosch et al., 2019; Neerincx et al., 2018; Nikolaidis et al., 2017).
@@
Adaptation in AI behavior, aimed at enhancing collaboration, involves learning from relevant human features and modifying behavior to align with human goals, thereby fostering a more intuitive and effective partnership (Michelle Zhao, 2022). Incorporating computational models of emotion into mental models allows AI agents to consider affective states in decision-making, enriching the collaboration with sophisticated, goal-oriented actions (Kaptein et al., 2016). Establishing a common vocabulary and shared understanding of goals, capabilities, and emotional states further solidifies the foundation of effective Human-AI collaboration (van den Bosch et al., 2019).
@@
Recent research initiatives showcase practical applications of this synergy. AdaTest++, an augmented tool for auditing large language models (LLMs), exemplifies the collaborative effort between humans and AI in enhancing model reliability and understanding, facilitating the identification of model failures through a combined effort (Rastogi et al., 2023). In the realm of mental health support, the HAILEY system demonstrates how AI can augment human empathy with analytical insights, providing feedback-driven support for more empathetic and effective peer-to-peer conversations (Sharma et al., 2023). Additionally, the innovative integration of LLMs and diffusion models for creating visual metaphors highlights the creative potential unleashed by human-AI collaboration, where human-proposed linguistic metaphors are visually realized through AI-generated imagery, bridging the gap between conceptual ideas and visual representation (Chakrabarty et al., 2023).
3.3. Compatibility of Human-AI Systems
@@ 3.3.1. Backwards Compatibility
In the realm of AI technologies, backward compatibility plays a pivotal role in ensuring the smooth integration of new or updated systems with pre-existing frameworks. This concept is crucial when embedding AI into legacy systems, as it involves careful consideration of existing tools, processes, and practices. Weisz et al. (Weisz et al., 2021) delve into how generative models contribute to the modernization of applications, emphasizing the necessity for AI systems to function cohesively within older infrastructures. In the context of traditional crowd work, Chhibber et al. (Chhibber et al., 2022) illustrate the significance of aligning AI tools with established workflows, showcasing the potential of generative models, like Generative Adversarial Networks (GANs) and autoencoders, to augment and refine datasets in legacy systems. Additionally, Bau et al. (Bau et al., 2019) present a novel approach using GANs to adapt image priors to specific image statistics in individual photographs, thereby tackling challenges in reproducing and manipulating high-level image attributes, as evidenced in various semantic editing tasks.
@@
Discussing the broader compatibility challenges outside just the technological perspective, Wang (Wang, 2023) talks about the transformative effects of Large Language Models (LLMs) on the job market, underscoring the necessity of integrating these advanced technologies into existing economic structures without causing major disruptions. Eloundou et al. (Eloundou et al., 2023) examine the impact of LLMs on the US labor market, revealing that approximately 80% of workers could experience a significant change in their tasks, with 19% potentially seeing over half of their tasks affected. This ‘exposure’ refers to a 50% increase in task efficiency without compromising quality. Furthermore, Harrer et al. (Harrer, 2023) highlight the potential of LLMs in generative AI applications, emphasizing the need for responsible design, human oversight, and ethical considerations to prevent the spread of misinformation. They offer insights into how these models, when developed responsibly, can serve as efficient and trustworthy tools in sectors like healthcare, addressing technology, risks, and ethical issues.
3.3.2. Learning Curve and Effective Usability
@@
In Human-AI Teaming (HAT), a critical factor is the ease and efficiency of interaction between humans and AI which significantly influences the learning curve and overall usability. McNeese et al.’s (McNeese et al., 2021) study talks about the team composition in HAT, revealing that AI’s ability to take a leading role in teams can enhance performance, though it often presents an initial learning curve for human participants. Lyons et al. (Lyons et al., 2019) discuss the importance of intuitive interfaces in fostering a partnership, rather than a tool-based relationship, between machines and humans. A pivotal aspect in this regard is the AI’s ability to explain its decisions, especially when those decisions may seem counter-intuitive to human collaborators. Hou et al.’s (Hou et al., 2023) research highlights that trust in AI is bolstered when humans understand the rationale behind AI decisions, fostering a sense of vulnerability and reliance on AI actions. This concept is further supported by Bansal et al. (Bansal et al., 2021), who find that AI explanations significantly increase human trust in the algorithm.
@@
Exploring the potential of Large Models, Wu et al. (Wu et al., 2022) also explore that breaking down the complex human tasks into multiple LLM prompts and chaining them together helps with increased human-AI interaction towards improvement. The users improvised over the new ways of interacting with the LLMs by leveraging the sub-tasks to set calibrate model expectations and making new-unit tests for each submodule of the LLM chain. Veselovsky et al.’s (Veselovsky et al., 2023) study, however, suggests that even though prevalent usage of LLMs generating high-quality but homogeneous responses might not truly capture the real essence of human behavior, they can prove to be useful too. They analyze that human workers not using LLMs in their tasks produced lower quality (less essential) information than when allowed the use of LLMs. They, hence, given the rapidly evolving large model landscape, emphasize a need to constantly monitor the co-evolution of LLM and human teaming to get to some constructive contributions.
3.3.3. Evaluating the Effectiveness of Joint Human-AI Systems